OpenStudy (janyia):

How to find the factor form of 28n^4+16n^3-80n^2=0

OpenStudy (amorfide):

you want to find the largest factor of n, in common with all terms we can start by taking out a factor of n at a time factor out n n(28n^2 +16n^2-80n)=0 but you can still take out an n since they are all still multiplied by n n(n(28n+16n-80))=0 you can't take any more factors of n out of these numbers, since -80 is not multiplied by any n so you have n(n)(28n+16n-80)=0 nxn=n^2 n^2(28n+16n-80)=0 now, find a number that goes into 28, 16, and -80. you want the largest number you can find I would start with 4 and factor out 4 n^2(4(7n^2+4n-20))=0 4n^2(7n^2 +4n -20))=0 now you want to factorise 7n^2 +4n -20 7n(n+2)-10(n+2) so your final answer will be 4n^2(7n-10)(n+2)

OpenStudy (amorfide):

that should be 4n^2(7n-10)(n+2)=0

OpenStudy (janyia):

Thanks so much